I need help with these 2 problems, and I beg you to please explain the equations step by step because I am just a FITH GRADER so I can understand it. Thanks a lot in advance.
1) Betty ate one third of her strawberries for snack. She had 12 left in her basket. How many did she have in the basket before she had her snack?
2) Sonia spent one third of her money, she then lost half of what she had left. She then had $ 6. How much money did she start with?



Answer by htmentor(1343)

1) Since she ate 1/3 of her strawberries, 2/3 of them remain.
We are told there are 12 left.
That means that 2/3 of the original amount, A, is 12:
2/3A = 12
So, the original amount is:
A = 3/2*12 = 18 strawberries
2) Let x = the amount of money she started with
If she spent 1/3 of her money, she has 2/3 left.
Then she lost half of what was left, which means she has half of the new amount remaining. The amount remaining is $6.
So, in equation form this is:
(1/2)*(2/3)x = 6.
Solve for x:
x = (3/2)*2*6 = $18